Differences

This shows you the differences between two versions of the page.

Link to this comparison view

Both sides previous revision Previous revision
operations:agenda2021_11_25 [2021/12/08 19:51]
Jasper Edwards
operations:agenda2021_11_25 [2021/12/08 19:55] (current)
Jasper Edwards
Line 1: Line 1:
-Chair:  Jasper+Chair: Pradyumna
  
 Next meeting chair: Lim Next meeting chair: Lim
  
-Attended:+Attended: Eloise, Simin, Jasper, Lim, Angira, David, Jamie
  
 Apologies: Apologies:
  
 General Issues: General Issues:
-*[25/11] rv150 - Yarragadee: mk5 crashed and was inaccessible, Had to call Yg to reboot the mk5 (Pradyumna). 
-*[03/12] r41027 - Yarragadee: Delayed start due to faulty autocorrelations (Pradyumna). 
-*[04/12] Si1337 - Hobart 12m:  fs kept freezing every time while setting up for exp.  Terminating and resetting didn't rectify the issue (Jay). 
-* [07/12] AUM043: 
-- Unable to VNC into ops8 … :1. Jay started second server (ops8 … :2), able to carry on with observations on this one (Eloise) 
- 
-Hobart 12m - maser not available? clkoff and maserdelay no reading, cannot start vnc maserhb, cannot read wth from MET sensor. upsmonhb computer not running, so counters not communicating. clkoff maserdelay now working. (Lim/warren) 
-Katherine 12m 
- 
-TCP/IP connection problem -> for above errors wake up comms every 10 minutes; antenna=open@!,10m, antenna=operate@!,10m , onsource@!,10m (warren) 
-Bad samplers -> reconfiguring dbbc3ke didn't work (Simin). 
- 
-“ERROR st -5 Error return from antenna, see Mbus error”, antenna=open/antenna=operate didn't work, halted schedule and reset drives on timeke. (Eloise) 
-Above error would happen every ~hour, antenna=open would work some of the time. 
  
 +Yg had a connection issues on two occasions (aov065 and r11025).
 +DBBC3 uses the abef samplers and the 'c' and 'd' displaying bad stats should not be a concern? ABEF are used for the S/X setting. Needs updating on the wiki.
 +2 new observers (David and Angira) to familiarize themselves this December. Ideally start their shifts in January. Change in roster needed. David Schuunk will be starting early next year as well.
 +First wideband vgos session held (vo1327). The session ran free of problems. Expected feedback on it from Haystack.
 +AOV065
  
 +There was a UPS failure which got fixed eventually after a manual power reboot (Ke).